### v2

@@ 2 Model and preliminaries
[…]
These are called the thin flow conditions (sometimes normalized thin flows or thin flows with resetting). We will take this viewpoint throughout, and always think of an equilibrium as a pair $(x,\ell)$ satisfying these conditions. Such a pair fully determines the flow over time $(f^{+},f^{-})$ for the equilibrium; for example, the queue volume is determined by $\tfrac{1}{\nu_{e}}z_{e}(\ell_{v}(\theta))=\max\{0,\ell_{w}(\theta)-\ell_{v}(\theta)-\tau_{e}\}$ for all $e=vw$ and $\theta\in\mathbb{R}_{\geq 0}$. Note that from this perspective, given $(x,\ell)$ satisfying the thin flow conditions for all times $\theta\in[0,T]$, extending the equilibrium to later departure times only requires knowing $\ell(T)$; nothing about the earlier (with respect to departure time) behavior of the equilibrium is needed. Just $\ell$ alone captures all the truly important information; one could easily verify whether a given curve $\ell$ can be extended to an equilibrium, since given $\ell^{\prime}(\theta)$, it is easy to determine whether a matching $x^{\prime}(\theta)$ exists so that $(x^{\prime}(\theta),\ell^{\prime}(\theta))$ satisfy the thin flow conditions. We call such an $\ell$ an equilibrium trajectory.
Given a pair $(x,\ell)$, we say that $\theta$ is a point of differentiability if both $x^{\prime}(\theta)$ and $\ell^{\prime}(\theta)$ exist. Almost every $\theta$ is a point of differentiability.
Note that the thin flow conditions (2) are fully determined by the pair $E^{\prime}$ and $E^{*}$. We call a pair $(E^{\prime},E^{*})$ with $E^{*}\subseteq E^{\prime}\subseteq E$ a valid configuration if (i) for every node $v\in V$ there is an $s$-$v$-path in $E^{\prime}$, (ii) every arc $e\in E^{*}$ lies on an $s$-$t$-path in $E^{\prime}$, and (iii) no arc of $E^{*}$ lies on a directed cycle in $(V,E^{\prime})$.
If $(E^{\prime},E^{*})$ is a valid configuration, then the thin flow conditions always have a solution [CCL15], and $\ell^{\prime}$ is uniquely determined [Koc12]. (Uniqueness in the more restrictive case that $E^{\prime}$ is acyclic was shown by [CCL15]; for most purposes this already suffices.)
Note that for any valid configuration $(E^{\prime},E^{*})$, the solution $(x^{\prime},\ell^{\prime})$ to the thin flow equations satisfy $\ell^{\prime}_{v}\leq\kappa$, where $\kappa\coloneqq u_{0}/\min_{e\in E}\nu_{e}$. Thus any equilibrium trajectory is $\kappa$-Lipschitz.
It has been shown [KS11, CCL15] that in an equilibrium, $\ell$ is a piecewise linear function; each linear segment is referred to as a phase of the equilibrium. However, it is not known whether the number of phases is necessarily finite, even within a compact interval.
We call a vector $l^{\lozenge}\in\mathbb{R}^{V}$ feasible if $(E^{\prime}_{l^{\lozenge}},E^{*}_{l^{\lozenge}})$ is a valid configuration, and define $\Omega\subseteq\mathbb{R}^{V}$ to be the subset of feasible labelings. Condition (i) for a valid configuration is simply insisting that $l^{\lozenge}$ does correspond to earliest arrival times from $s$, for some choice of nonnegative queue lengths on the arcs. Condition (ii) ensures that a flow particle at the back of a queue travels on an active $s$-$t$-path, which is the definition of dynamic equilibria. Condition (iii) is always satisfied, for any $l^{\lozenge}\in\mathbb{R}^{V}$, because $E^{\prime}_{l^{\lozenge}}$ is always acyclic: we have that $\tau_{e}\geq 0$ for all $e\in E$ and moreover, that the sum of transit times along a directed cycle is strictly positive. Thus, no cycle is completely contained in shortest paths from $s$ [^1] We do not really need that the sum of transit times along a cycle is strictly positive; we could observe the same by exploiting that it takes strictly positive time to traverse an arc with a queue.
@@ 3 Technical overview
Equilibrium trajectories.
We first further develop the view of equilibria as trajectories whose derivatives are controlled by a vector field.
Let $X:\Omega\to\mathbb{R}^{V}$ be the vector field for which $X(l^{\circ})$ is the solution to the thin flow equation for $(E^{\prime}_{l^{\circ}},E^{*}_{l^{\circ}})$, for all $l^{\circ}\in\Omega$. So if $\ell$ is an equilibrium trajectory, $\ell^{\prime}(\theta)=X(\ell(\theta))$ for almost every $\theta$. Now since $X(l^{\circ})$ depends only on $E^{\prime}_{l^{\circ}}$ and $E^{*}_{l^{\circ}}$, we can deduce that $X$ is piecewise constant. The regions on which $X$ is constant also have a very simple structure, given by (1): each arc $e=vw$ divides $\Omega$ into two open halfspaces separated by the hyperplane $\{l^{\circ}\in\Omega:l^{\circ}_{w}-l^{\circ}_{v}=\tau_{e}\}$; a region is determined by the choice of sign (positive, negative or zero) for each link (not all combinations necessarily yield a region); see Figure 4.
@@
A more difficult situation is shown in Figure 4, where the trajectory (informally) spirals outwards. Again, non-uniqueness is rather clear; two possible trajectories starting from the origin are shown. This possibility is not excluded by previous work. In particular, if a trajectory of this type existed, it would provide an example of an equilibrium $(x,\ell)$ for which $\ell$ is not right-differentiable. By contrast, the procedure of [CCL15] and Koch and Skutella [KS11] produces an equilibrium whose labeling is right-differentiable, and so this would immediately be an example of non-uniqueness.
@@
For any $l^{\circ}\in\Omega$, there is a unique equilibrium trajectory $\ell$ with $\ell(0)=l^{\circ}$.
@@
Let $\mathcal{L}$ denote the Banach space of $\kappa$-Lipschitz functions from $[0,\infty)$ to $\mathbb{R}$ imbued with the uniform norm $\|g\|_{\infty}=\sup_{\theta\geq 0}g(\theta)$. (Recall that $\kappa=u_{0}/\min_{e}\nu_{e}$ is a Lipschitz constant for all equilibrium trajectories.) By $\mathcal{X}$ we denote the normed space $\mathcal{L}^{V}$ imbued with the norm $\|\ell\|\coloneqq\max_{v\in V}\|\ell_{v}\|_{\infty}$. Our continuity result is the following.
@@
Let $\Psi:\Omega\to\mathcal{X}$ be the map that takes $l^{\circ}\in\Omega$ to the unique equilibrium trajectory $\ell$ satisfying $\ell(0)=l^{\circ}$. Then $\Psi$ is a continuous map.
We remark that the continuity of $\Psi$ is a statement over the entire evolution of a trajectory—a rather strong form of continuity, and in particular, stronger than convergence of the trajectory at any fixed time $\theta$. Our result says that if we look at the equilibrium trajectory $\ell$ with $\ell(0)=l^{\circ}$, and then look at the equilibrium trajectory $\tilde{\ell}$ with ~$\tilde{\ell}(0)$ equal to a slight perturbation of $l^{\circ}$, then $\ell$ and $\tilde{\ell}$ stay close forever, rather than possibly drifting apart very slowly.
@@
To understand the connection between long-term behavior and uniqueness and continuity, we consider the “local” situation in a sufficiently small neighborhood of some $l^{\circ}\in\Omega$. All hyperplanes that do not go through this point are considered to be very far away, and hence are ignored. To formalize this idea, we introduce generalized networks.
A generalized network is a graph $\hat{G}=(V,\hat{E})$, together with a set of special arcs $E^{\infty}$ such that $(\hat{E},E^{\infty})$ is a valid configuration. We consider all arcs in $E^{\infty}$ as always resetting; this can be interpreted as allowing the queues on these arcs to become negative, and thus we call them free arcs. The notion of a valid configuration and hence feasible labeling should be adjusted in the obvious way for this local instance: $(\hat{E}^{\prime},\hat{E}^{*})$ is a valid configuration if $E^{\infty}\subseteq\hat{E}^{*}\subseteq\hat{E}^{\prime}\subseteq\hat{E}$ and properties (i) to (iii) are fulfilled. The feasible set thus changes; denote it by $\hat{\Omega}$.
Given a specific point $l^{\circ}$, the relevant generalized network to consider, which we call the local network is the following. Call an arc $e=vw$ tight (with respect to $l^{\circ}$) if $l^{\circ}_{w}-l^{\circ}_{v}=\tau_{e}$; so the hyperplane of arc $e$ passes through $l^{\circ}$ precisely if it is tight. All non-tight arcs either have a queue or are inactive. All the arcs with a queue become free arcs $E^{\infty}$, and all inactive arcs are deleted. In other words, $\hat{E}=E_{l^{\circ}}^{\prime}$ and $E^{\infty}=E^{*}_{l^{\circ}}$. Only arcs of $\hat{E}\setminus E^{\infty}$ have an associated hyperplane; all these hyperplanes pass through $l^{\circ}$. In other words, the vector field $X$ of the local network has a conic self-similarity: $X(l^{\circ}+\alpha d)=X(l^{\circ}+d)$ for any $d\in\mathbb{R}^{V}$ with $l^{\circ}+d\in\hat{\Omega}$ and $\alpha>0$.
@@
We are now ready to make the connection between uniqueness/continuity and long-term behavior of equilibrium trajectories in this conic setting. Let $(y,\lambda)$ be the solution of the thin flow problem at $l^{\circ}$; so $\lambda=X(l^{\circ})$. Suppose that we could show that for some given initial condition $\hat{l}^{\circ}$, any equilibrium trajectory $\ell$ starting from $\hat{l}^{\circ}$ eventually, after some finite amount of time $T$, satisfies $\ell^{\prime}(\theta)=\lambda$ for all $\theta\geq T$. Here, $T$ may certainly depend on the choice of $\hat{l}^{\circ}$. But now we can exploit the conic symmetry. Consider an equilibrium trajectory $\ell^{(\varepsilon)}$ starting from $l^{\circ}+\varepsilon(\hat{l}^{\circ}-l^{\circ})$. Then it is quite easy to see that the trajectory $\ell$ defined by $\ell(\theta)\coloneqq(\ell^{(\varepsilon)}(\varepsilon\theta)-l^{\circ})/\varepsilon+l^{\circ}$ is an equilibrium trajectory starting from $\hat{l}^{\circ}$. Thus $\ell^{(\varepsilon)}(\theta)$ will reach steady state at time $\varepsilon T$. So as we move the initial condition closer and closer to $l^{\circ}$, any equilibrium trajectory looks more and more like the trajectory $\ell^{*}$ given by $\ell^{*}(\theta)=l^{\circ}+\theta\lambda$. The maximum distance between $\ell^{(\varepsilon)}(\varepsilon)$ and $\ell^{*}(\varepsilon)$ can be controlled by the distance between the initial condition and $l^{\circ}$ and by exploiting Lipschitz continuity of the trajectories. This shows uniqueness at $l^{\circ}$. Let $\ell$ be any equilibrium trajectory starting from $l^{\circ}$. For any $\varepsilon>0$, we can think of $\ell$ on $[\varepsilon,\infty)$ as an equilibrium trajectory starting from $\ell(\varepsilon)$. Since $\ell(\varepsilon)\to l^{\circ}$ as $\varepsilon\to 0$, it follows from the above that $\sup_{\theta\geq 0}[\ell(\theta+\varepsilon)-\ell^{*}(\theta)]$ converges to 00 as $\varepsilon\to 0$. So $\ell=\ell^{*}$.
@@
Note that while this conic view gives some nice intuition, our formal proof does not rely on this explicitly. Instead, we show directly that not only does every trajectory reaches steady state in some finite time, but that the time required to reach steady state scales with the distance between the initial point and $l^{\circ}$.
@@
This also shows continuity locally around $l^{\circ}$: equilibrium trajectories starting from small perturbations of $l^{\circ}$ remain close to the equilibrium trajectory starting from $l^{\circ}$. Essentially, this shows continuity of the trajectories over a single phase, and it is not too difficult to deduce continuity over the entire equilibrium from this.
@@
Let $G=(V,E)$ and $E^{\infty}\subseteq E$ be a generalized network, $l^{\circ}$ any feasible labeling, and $\lambda$ the solution to the thin flow equations for configuration $(E,E^{\infty})$. Then for any equilibrium trajectory $\ell$ with $\ell(0)=l^{\circ}$, there exists some $T$ such that $\ell^{\prime}_{v}(\theta)=\lambda_{v}$ for all $v\in V$ and $\theta\geq T$.
One can give explicit bounds on $T$ in terms of the instance and $l^{\circ}$; we do so in Section 4. Note that in the theorem we show that after time $T$ all labels change as in the thin flow $\lambda$. This is a stronger notion of steady state than asking only that queues change with the right rates as described in Section 1. We will use the stronger notion as the definition of steady state.
For uniqueness and continuity, the theorem is applied with $E^{\infty}=E^{*}_{l^{\circ}}$. The choice $E^{\infty}=\emptyset$ is of independent interest: it corresponds to a real equilibrium trajectory with no artificial free arcs. As mentioned earlier, [CCO21] proved the following:
@@
Let us first summarize the approach taken by [CCO21] under the inflow condition and $E^{\infty}=\emptyset$. They first pose and answer the following question (paraphrased): which choices of initial condition $l^{\circ}$ have the property that $\ell(\theta)=l^{\circ}+\lambda\theta$ is immediately an equilibrium trajectory ? (In their case, the thin flow direction $\lambda$ is the all-ones vector.) They show that the answer is provided in full by considering the following primal-dual LP:
@@
Dual optima $(d^{*},p^{*})$ are in one-to-one correspondence with steady-state initial conditions: $d^{*}$ represent a possible initial labeling, and $p^{*}_{e}$ the queue length on arc $e$. Primal optima, on the other hand, are in one-to-one correspondence with equilibria departure flows. In other words: $(x,\ell)$ in which $\ell_{v}(\theta)=d^{*}_{v}+\theta$ for some dual optimum $d^{*}$, and $x^{\prime}(\theta)$ is a primal optimum for almost every $\theta$, is an equilibrium; and all equilibria that are in steady state from the very beginning, are of this form.
It is by no means immediately apparent why answering this question regarding the characterization of steady states is helpful in proving convergence to steady state. The key novelty in [CCO21] is that the dual LP provides us with the correct potential function. Namely, they define, given an equilibrium $(x,\ell)$ with corresponding queueing delays $q$, the potential function
@@
It turns out that while a primal-dual LP characterizing the steady state is still the key to producing the correct potential, the situation is much more subtle. The obvious generalization of (3), with $\nu_{e}$ replaced by $\hat{\nu}_{e}$ and no other changes, is not the correct one. Rather, one must observe that there is a larger class of candidate LPs, from which a choice must be carefully made. Let $y\in\mathbb{R}_{\geq 0}^{V}$ be such that $(y,\lambda)$ is a thin flow for configuration $(E,\emptyset)$. A first observation is that for arcs $e=vw$ with $\lambda_{w}>\lambda_{v}$, we may enforce the constraint $f_{e}=\hat{\nu}_{e}$, and for arcs $e=vw$ with $\lambda_{w}<\lambda_{v}$, we may enforce the constraint $f_{e}=0$, without changing the feasible set. This comes from observing that if one looks at any set $S$ of the form $S=\{v:\lambda_{v}\leq t\}$, for some $t\in\mathbb{R}_{\geq 0}$, and this set is nontrivial (neither empty nor $V$), then $y_{e}=\lambda_{w}\nu_{e}$ for all $e=vw\in\delta^{+}(S)$ and $y_{e}=0$ for all $e\in\delta^{-}(S)$ by the thin flow conditions.
@@
Consider some instance, and the corresponding equilibrium trajectory $\ell$ starting (for simplicity) from the empty network at time 00. Now suppose we perturb the transit times slightly, leading to a new equilibrium trajectory $\hat{\ell}$ on the perturbed instance. We show that $\hat{\ell}$ remains similar to $\ell$.
@@
Similarly, we can consider perturbing the capacities $\nu_{e}$ and/or the inflow rate $u_{0}$. Again we get continuity, though necessarily of a slightly weaker form: we must restrict ourselves to a compact interval. This is because the steady state direction $\lambda$ for a phase does change as $\nu$ or $u_{0}$ changes—but only continuously. This means that once the final steady state of the overall equilibrium is reached, the trajectories may slowly diverge.

### v3

@@ 2.2 Equilibria
Assume for contradiction that $V_{0}\neq\emptyset$. As we choose a maximal thin flow solution, there has to be a resetting arc $f=yz$ with $x^{\prime}_{f}=0$. Otherwise, $\ell^{\prime}$ is not pointwise maximal: with $\varepsilon$ equal to the smallest nonzero label of $\ell^{\prime}$, $\bar{\ell}^{\prime}$ defined by $\bar{\ell}^{\prime}_{v}=\max\{\ell^{\prime}_{v},\varepsilon\}$ for all $v$ is easily seen to still be a solution to the thin flow equations.
Note that $z\in V_{0}$. As $(E^{\prime},E^{*})$ is a valid configuration, there has to be an active $z$-$t$-path. In particular there has to be an active arc $e=vw\in E^{\prime}$ with $v\in V_{0}$ and $w\notin V_{0}$. But since $x^{\prime}_{e}=0$, this is a contradiction to (TF-3).
The second part of the claim is an immediate consequence of (2); $x^{\prime}_{e}=0$ for $e=vw\in E^{*}$ would imply $\ell^{\prime}_{w}=0$. ∎
2.3 Equilibrium trajectories and generalized networks
The exposition so far demonstrates the primacy of the earliest arrival labels. If $X(l)$ is the thin flow direction for configuration $(E^{\prime}_{l},E^{*}_{l})$ wherever this is valid, then an equilibrium $(x,\ell)$ satisfies that $\ell^{\prime}(\theta)=X(\ell(\theta))$ for almost all $\theta\in\mathbb{R}_{\geq 0}$. This is the primary perspective on equilibria that we will take going forward. However, we will need to generalize it in two ways, for the purposes of our later arguments.
Firstly, we will allow for an arbitrary starting point, not necessarily the labeling corresponding to the empty network. Secondly, we allow for a specified set of “free arcs” $E^{\infty}\subseteq E$ to be given which will always be resetting (and hence also always active). In some sense, these represent arcs where the queue is allowed to be negative. The reason that this will be useful will become clearer later; for now, we focus on the precise definitions.
Definition 2.5.
A generalized network is a network $G=(V,E)$, along with a set $E^{\infty}\subseteq E$ of free arcs such that $(E,E^{\infty})$ is a valid configuration.
In particular, $E^{\infty}=\emptyset$ is always a valid choice. Fix a generalized network $G=(V,E)$ and $E^{\infty}$ for the remainder of this section.
We redefine $E^{\prime}_{l}$ and $E^{*}_{l}$ for $l\in\mathbb{R}^{V}$ by
$E^{\prime}_{l}=E^{\infty}\cup\{e=vw\in E\mid l_{w}-l_{v}\geq\tau_{e}\}\qquad\text{and}\qquad E^{*}_{l}=E^{\infty}\cup\{e=vw\in E\mid l_{w}-l_{v}>\tau_{e}\}.$ (2)
As before, $l$ is a feasible labeling if $(E^{\prime}_{l},E^{*}_{l})$ is a valid configuration. Let $\Omega$ denote the set of feasible labelings.
We remark that $E^{\prime}_{l}$ is acyclic for any $l\in\Omega$, just as was the case for normal networks. This follows as the set $\{e=vw\in E\mid l_{w}-l_{v}-\tau_{e}\geq 0\}$ is acyclic by our assumption that sum of transit times along a directed cycle is positive, and moreover, no arc $E^{\infty}$ lies on a cycle of active arcs as $(E_{l}^{\prime},E_{l}^{*})$ is a valid configuration.
The following lemma shows that a trajectory that follows the vector field $X$ will not leave $\Omega$. (The existence results in the literature certainly imply this in the case that $E^{\infty}=\emptyset$ and $\ell(0)$ is the labeling of an empty network.)
Lemma 2.6.
For every $l\in\Omega$, $X(l)\in\{d\in\mathbb{R}^{V}:l+\varepsilon d\in\Omega\text{ for some }\varepsilon>0\}$.
Proof.
Fix $l\in\Omega$, and let $l^{\prime}=X(l)$. Let $\varepsilon>0$ be small enough that with $\bar{l}=l+\varepsilon l^{\prime}$, for any $e=vw$ with $l_{w}-l_{v}-\tau_{e}\neq 0$, $\bar{l}_{w}-\bar{l}_{v}-\tau_{e}\neq 0$ and has the same sign. We need to show that properties (i)–(iii) for a valid configuration hold for $(E^{\prime}_{\bar{l}},E^{*}_{\bar{l}})$. Note that $E^{*}_{\bar{l}}\supseteq E^{*}_{l}$ and $E^{\prime}_{\bar{l}}\subseteq E^{\prime}_{l}$.
  1. (i)
    Suppose for a contradiction that there is no $s$-$v$ path for some $v\in V$, chosen minimally in the topological order described by $E^{\prime}_{\bar{l}}$. Then no arc entering $v$ is in $E^{\prime}_{\bar{l}}$, and hence in $E^{*}_{l}$. By (TF-3) (which holds since $E^{\prime}_{l}$ is acyclic), and as no arcs entering $v$ are in $E^{*}_{l}$, there exists some $uv\in E^{\prime}_{l}$ with $l^{\prime}_{v}=l^{\prime}_{u}$. But then $\bar{l}_{v}-\bar{l}_{u}-\tau_{e}=l_{v}-l_{u}-\tau_{e}\geq 0$, contradicting out assumption that $uv\notin E^{\prime}_{\bar{l}}$.
  2. (ii)
    Let $e=vw\in E^{*}_{\bar{l}}$. Take $(l^{\prime},x^{\prime})$ to be any complete solution to the thin flow equations. Either $l^{\prime}_{w}>l^{\prime}_{v}$, or $e\in E^{*}_{l}$. Either way (with the help of Section 2.2 in the latter case), $x^{\prime}_{e}>0$.
    Let $P$ be an $s$-$t$-path including $e$ in the support of $x^{\prime}$. The thin flow equations imply that each $f=yz\in P$ is either in $E^{*}_{l}$ (and hence $E^{*}_{\bar{l}}$), or is active and has $l^{\prime}_{z}\geq l^{\prime}_{y}$; so this path is in $E^{\prime}_{\bar{l}}$.
  3. (iii)
    Since $E^{\prime}_{\bar{l}}\subseteq E^{\prime}_{l}$ and $E^{\prime}_{l}$ is acylic, this is immediate.
Definition 2.7.
An equilibrium trajectory $\ell$ is a vector $(\ell_{v})_{v\in V}$, with $\ell_{v}:\mathbb{R}_{\geq 0}\to\mathbb{R}^{V}$ differentiable almost everywhere, such that $\ell(\theta)\in\Omega$ and $\ell^{\prime}(\theta)=X(\ell(\theta))$ for all $\theta\in\mathbb{R}_{\geq 0}$. The starting point of $\ell$ is the labeling $\ell(0)$.
The previous lemma implies that for any $l^{\circ}\in\Omega$, an equilibrium trajectory starting from $l^{\circ}$ can be constructed, by solving the differential equation $\ell(0)=l^{\circ}$, $\ell^{\prime}(\theta)=X(\ell(\theta))$ for $\theta\in\mathbb{R}_{\geq 0}$, exactly as was done for equilibria where $l^{\circ}$ is the labeling for the empty network.
Notice that since $X(l^{\circ})$ depends only on $E^{\prime}_{l}$ and $E^{*}_{l}$, we can deduce that $X$ is piecewise constant. The regions on which $X$ is constant also have a very simple structure, given by (2): each arc $e=vw$ divides $\Omega$ into two open halfspaces separated by the hyperplane $\{l\in\Omega:l_{w}-l_{v}=\tau_{e}\}$; a region is determined by the choice of sign (positive, negative or zero) for each link (not all combinations necessarily yield a region); see Figure 4.
Intuitively, an equilibrium trajectory starting from $l^{\circ}$ corresponds to an equilibrium, but with some initial conditions defined by $l^{\circ}$. However, interpreting this in terms of artificial initial queues involves some subtleties. We emphasize that no such interpretation is needed; equilibrium trajectories with initial conditions not corresponding to an empty network, or with free arcs, are tools for our analysis.
Lemma 2.8.
Any equilibrium trajectory is $\kappa$-Lipschitz, where
$\kappa\coloneqq\max\Bigl\{1,u_{0}/\min_{e\in E}\nu_{e}\Bigr\}.$
Proof.
Consider any equilibrium trajectory $\ell$. It suffices to show that for any point of differentiability $\theta$, $\ell^{\prime}_{v}(\theta)\leq\kappa$ for all $v\in V$. Let $(\ell^{\prime},x^{\prime})$ be a complete maximal thin flow solution for the configuration $(E^{\prime}_{\ell(\theta)},E^{*}_{\ell(\theta)})$ (so $\ell^{\prime}=\ell^{\prime}(\theta)$). Then for any $e=vw\in E^{\prime}_{\ell(\theta)}$,
$\ell^{\prime}_{w}\leq\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})\leq\max\{\ell^{\prime}_{v},x^{\prime}_{e}/\nu_{e}\}\leq\max\{\ell^{\prime}_{v},u_{0}/\min_{e^{\prime}\in E}\nu_{e^{\prime}}\},$
where the last inequality uses that $x^{\prime}_{e}\leq u_{0}$ by the acyclicity of $E^{\prime}_{\ell(\theta)}$. Considering any active $s$-$v$-path immediately demonstrates the claim.
@@ 3 Technical overview
A more difficult situation is shown in Figure 4, where the trajectory (informally) spirals outwards. Again, non-uniqueness is rather clear; two possible trajectories starting from the origin are shown. This possibility is not excluded by previous work. In particular, if a trajectory of this type existed, it would provide an example of an equilibrium $(\ell,x)$ for which $\ell$ is not right-differentiable. By contrast, the procedure of Cominetti, Correa and Larré [CCL15] and Koch and Skutella [KS11] produces an equilibrium whose labeling is right-differentiable, and so this would immediately be an example of non-uniqueness.
@@
Given a generalized network $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$ and a feasible labeling $l^{\circ}$, there is a unique equilibrium trajectory $\ell$ starting from $l^{\circ}$.
This theorem with $E^{\infty}=\emptyset$ and $l^{\circ}$ the labeling of the empty network gives label-uniqueness of equilibria.
@@
Let $\mathcal{L}$ denote the extended Banach space of $\kappa$-Lipschitz functions from $0,\infty)$ to $\mathbb{R}$ imbued with the extended [^3] An extended norm is allowed to take on infinite values, and otherwise behaves as a norm. norm $\|g\|_{\infty}=\sup_{\theta\geq 0}g(\theta)$. (Recall that $\kappa$ is a Lipschitz constant for all equilibrium trajectories by Section 2.3.) By $\mathcal{X}$ we denote the space $\mathcal{L}^{V}$ imbued with the extended norm $\|\ell\|\coloneqq\max_{v\in V}\|\ell_{v}\|_{\infty}$. Our continuity result is the following.
@@
Consider a generalized network $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$ and a feasible labeling $l^{\circ}$. Let $\Psi:\Omega\to\mathcal{X}$ be the map that takes $l^{\circ}\in\Omega$ to the unique equilibrium trajectory $\ell$ starting from $l^{\circ}$. Then $\Psi$ is a continuous map.
We remark that the continuity of $\Psi$ is a statement over the entire evolution of a trajectory — a rather strong form of continuity, and in particular, stronger than convergence of the trajectory at any fixed time $\theta$. Our result says that if we look at the equilibrium trajectory $\ell$ starting from $l^{\circ}$, and then look at the equilibrium trajectory $\tilde{\ell}$ starting from a slight perturbation of $l^{\circ}$, then $\ell$ and $\tilde{\ell}$ stay close forever, rather than possibly drifting apart very slowly.
@@
[…]
@@
We are now ready to make the connection between uniqueness/continuity and long-term behavior of equilibrium trajectories in this conic setting. Let $(y,\lambda)$ be the solution of the thin flow problem at $l^{\circ}$; so $\lambda=X(l^{\circ})$. Suppose that we could show that for some given initial condition $^{\color[rgb]{0,0,0}\hat{l}}$, any equilibrium trajectory $\ell$ starting from $^{\color[rgb]{0,0,0}\hat{l}}$ eventually, after some finite amount of time $T$, satisfies $\ell^{\prime}(\theta)=\lambda$ for all $\theta\geq T$. Here, $T$ may certainly depend on the choice of $^{\color[rgb]{0,0,0}\hat{l}}$. But now we can exploit the conic symmetry. Consider an equilibrium trajectory $\ell^{(\varepsilon)}$ starting from $l^{\circ}+\varepsilon({\color[rgb]{0,0,0}\hat{l}}-l^{\circ})$. Then it is quite easy to see that the trajectory $\ell$ defined by $\ell(\theta)\coloneqq(\ell^{(\varepsilon)}(\varepsilon\theta)-l^{\circ})/\varepsilon+l^{\circ}$ is an equilibrium trajectory starting from $^{\color[rgb]{0,0,0}\hat{l}}$. Thus $\ell^{(\varepsilon)}(\theta)$ will reach steady state at time $\varepsilon T$. So as we move the initial condition closer and closer to $l^{\circ}$, any equilibrium trajectory looks more and more like the trajectory $\ell^{*}$ given by $\ell^{*}(\theta)=l^{\circ}+\theta\lambda$. The maximum distance between $\ell^{(\varepsilon)}(\varepsilon)$ and $\ell^{*}(\varepsilon)$ can be controlled by the distance between the initial condition and $l^{\circ}$ and by exploiting Lipschitz continuity of the trajectories, which holds by Section 2.3. This shows uniqueness at $l^{\circ}$. Let $\ell$ be any equilibrium trajectory starting from $l^{\circ}$. For any $\varepsilon>0$, we can think of $\ell$ on $\varepsilon,\infty)$ as an equilibrium trajectory starting from $\ell(\varepsilon)$. Since $\ell(\varepsilon)\to l^{\circ}$ as $\varepsilon\to 0$, it follows from the above that $\sup_{\theta\geq 0}[\ell(\theta+\varepsilon)-\ell^{*}(\theta)]$ converges to 0 as $\varepsilon\to 0$. So $\ell=\ell^{*}$.
@@
Note that while this conic view gives some nice intuition, our formal proof does not rely on this explicitly. Instead, we show directly that not only does every trajectory reach steady state in some finite time, but that the time required to reach steady state scales with the distance between the initial point and $l^{\circ}$.
@@
This also shows continuity locally around $l^{\circ}$: equilibrium trajectories starting from small perturbations of $l^{\circ}$ remain close to the equilibrium trajectory starting from $l^{\circ}$. Essentially, this shows continuity of the trajectories over a single phase, and it is not too difficult to deduce continuity of the entire equilibrium trajectory from this.
@@
Consider a generalized network $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$. Let $\lambda$ be the thin flow direction for configuration $(E,E^{\infty})$, and let $l^{\circ}$ be any feasible labeling. Then there is a value $T$ such that for any equilibrium trajectory $\ell$ starting from $l^{\circ}$, $\ell^{\prime}_{v}(\theta)=\lambda_{v}$ for all $v\in V$ and $\theta\geq T$.
One can give explicit bounds on $T$ in terms of the instance and $l^{\circ}$; we do so in Section 4. Note that in the theorem we show that after time $T$ all labels change as per the thin flow direction $\lambda$. This is a stronger notion of steady state than asking only that queues change with the right rates as described in Section 1. We will use the stronger notion as the definition of steady state. We will also refer to $\lambda$ as the steady state direction.
For uniqueness and continuity, the theorem is applied to the current shortest path network, with $E^{\infty}=E^{*}_{l^{\circ}}$. The choice $E^{\infty}=\emptyset$ is of independent interest: it corresponds to a real equilibrium trajectory on a given instance with no artificial free arcs. As mentioned earlier, [CCO21] proved the following:
@@
Let us first summarize the approach taken by [CCO21] under the inflow condition and $E^{\infty}=\emptyset$. They first pose and answer the following question (paraphrased): which choices of initial condition $l^{\circ}$ have the property that $\ell(\theta)=l^{\circ}+\lambda\theta$ is immediately an equilibrium trajectory? (In their case, the steady-state direction $\lambda$ is the all-ones vector.) They show that the answer is provided in full by considering the following primal-dual LP:
@@
Dual optima $(d^{*},p^{*})$ are in one-to-one correspondence with steady-state initial conditions: $d^{*}$ represent a possible initial labeling, and $p^{*}_{e}$ the queue length on arc $e$. Primal optima, on the other hand, are in one-to-one correspondence with equilibria departure flows. In other words: $(\ell,x)$ in which $\ell_{v}(\theta)=d^{*}_{v}+\theta$ for some dual optimum $d^{*}$, and $x^{\prime}(\theta)$ is a primal optimum for almost every $\theta$, is an equilibrium; and all equilibria that are in steady state from the very beginning, are of this form.
It is by no means immediately apparent why answering this question regarding the characterization of steady states is helpful in proving convergence to steady state. The key novelty in [CCO21] is that the dual LP provides us with the correct potential function. Namely, they define, given an equilibrium $(\ell,x)$ with corresponding queueing delays $q$, the potential function
@@
It turns out that while a primal-dual LP characterizing the steady state is still the key to producing the correct potential, the situation is much more subtle. The obvious generalization of (3), with $\nu_{e}$ replaced by $\hat{\nu}_{e}$ and no other changes, is not the correct one. Rather, one must observe that there is a larger class of candidate LPs, from which a choice must be carefully made. Let $y\in\mathbb{R}_{\geq 0}^{{\color[rgb]{0,0,0}E}}$ be such that $(y,\lambda)$ is a thin flow for configuration $(E,\emptyset)$. A first observation is that for arcs $e=vw$ with $\lambda_{w}>\lambda_{v}$, we may enforce the constraint $f_{e}=\hat{\nu}_{e}$, and for arcs $e=vw$ with $\lambda_{w}<\lambda_{v}$, we may enforce the constraint $f_{e}=0$, without changing the feasible set. This comes from observing that if one looks at any set $S$ of the form $S=\{v:\lambda_{v}\leq t\}$, for some $t\in\mathbb{R}_{\geq 0}$, and this set is nontrivial (neither empty nor $V$), then $y_{e}=\lambda_{w}\nu_{e}$ for all $e=vw\in\delta^{+}(S)$ and $y_{e}=0$ for all $e\in\delta^{-}(S)$ by the thin flow conditions.
@@
Consider some instance, and the corresponding equilibrium trajectory $\ell$ starting (for simplicity) from the empty network at time 0. Now suppose we perturb the transit times slightly, leading to a new equilibrium trajectory $\hat{\ell}$ on the perturbed instance. We show that $\hat{\ell}$ remains similar to $\ell$.
@@
Similarly, we can consider perturbing the capacities $\nu_{e}$ and/or the inflow rate $u_{0}$. Again we get continuity, though necessarily of a slightly weaker form: we must restrict ourselves to a compact interval. This is because the steady state direction $\lambda$ does change as $\nu$ or $u_{0}$ changes — but only continuously. This means that once the final steady state of the overall equilibrium is reached, the trajectories may slowly diverge.